# Conditional Autoencoder for Generating BNS Waveforms with Tidal and Precession Effects

### v1

@@ VI Summary and Discussion
This study presents an efficient gravitational waveform generation method based on a cAE and applies it to amplitude-phase modeling of BNS systems. Compared to traditional waveform approximation methods such as IMRPhenomXP_NRTidalv2, cAE significantly improves computational efficiency while maintaining high reconstruction accuracy. On a large-scale test dataset, the averaged waveform mismatch is $3.97\times 10^{-3}$, corresponding to an average accuracy exceeding $99.6\%$. Even with precession and tidal effects, cAE maintains high precision across different $\Theta$ ranges, particularly in the high-cycle orbital evolution phase. In terms of efficiency, the cAE model demonstrates a significant advantage over traditional waveform approximation methods. For a single waveform, cAE requires only $0.12$ seconds, compared to $0.38$ seconds for IMRPhenomXP_NRTidalv2 and $0.62$ seconds for IMRPhenomPv2_NRTidal, yielding speedups of approximately $3$ to $5$ times. As the batch size increases, this advantage becomes more pronounced. For 1,000 waveforms, cAE completes the task in $0.86$ seconds, while IMRPhenomXP_NRTidalv2 and IMRPhenomPv2_NRTidal require $46.12$ and $53.47$ seconds, respectively—corresponding to speedups of around $54$ and $62$ times. When generating 10,000 waveforms, cAE requires only $7.48$ seconds, compared to $454.51$ seconds for IMRPhenomXP_NRTidalv2 and $497.75$ seconds for IMRPhenomPv2_NRTidal, achieving speedups of over $60$ times. These results suggest that cAE scales more efficiently with batch size compared to traditional methods. These results suggest that our cAE model is more efficient than traditional waveform approximation methods for generating waveforms at large batch size.
Additionally, this study constructs a large-scale BNS waveform dataset covering an extended physical parameter space, including mass, spin, and tidal deformation, allowing the model to learn waveform characteristics under complex physical conditions. To improve time-series modeling, the cAE architecture integrates ResNet and Transformer mechanisms, enhancing its ability to represent long-duration waveforms and reducing reconstruction errors in high-cycle orbital phases. This design improves generalization and offers new insights for high-precision waveform modeling.
Despite the progress achieved in this study, several aspects require further optimization. The analysis indicates a slight increase in mismatch for both high- and low-mass binaries. In high-mass systems, the short merger time limits the network’s ability to capture waveform details, while the relatively small number of training samples may constrain generalization. Data augmentation or adaptive sampling could mitigate these effects. For low-mass binaries, the higher number of orbital cycles presents additional challenges for accurate waveform fitting, suggesting room for improvement in modeling long-term evolution. The model’s performance under different spin and tidal deformation parameters shows less variation than expected, indicating that sensitivity to extreme parameter regions warrants further investigation. Future work will systematically analyze key factors influencing waveform generation to enhance physical consistency and generalization. While cAE significantly improves computational efficiency over traditional methods, further acceleration strategies, such as TensorRT[59, 60] or ONNX Runtime[61], could further reduce inference time, improving suitability for real-time GW analysis.
Several directions for future research emerge from this study. The impact of latent variable dimensionality on waveform accuracy and efficiency remains unexplored and will be systematically analyzed. Expanding the parameter space to include more complex sources, such as eccentric binaries and a broader range of spin and tidal effects, will further improve applicability. On the architectural side, advanced deep learning methods, including diffusion models [62] and Transformer variants [63], could enhance waveform generation accuracy and efficiency. Moreover, integrating deep learning with JAX-based acceleration techniques [64] may further improve computational efficiency. Additionally, Bayesian inference with neural networks for waveform generation could accelerate parameter estimation [65] for GW events.
@@ Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203004), the National Natural Science Foundation of China (Grant No. 12347101), and the Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-MSX0103).
@@
[…]
Figure 15: Corner plot of $3\times 10^{5}$ test set samples, showing the joint distribution of the 10-dimensional parameter space. The diagonal subplots display 1D histograms of each parameter, while the off-diagonal subplots show 2D joint distributions, where the color scale represents the corresponding mismatch value of each sample.

### v4

@@ VI Summary and Discussion
This study presents an efficient gravitational waveform generation method based on a cAE and applies it to amplitude-phase modeling of BNS systems. Compared to traditional waveform approximation methods such as IMRPhenomXP_NRTidalv2, cAE significantly improves computational efficiency while maintaining high reconstruction accuracy. On a large-scale test dataset, the averaged waveform mismatch is $2.13\times 10^{-3}$, corresponding to an average accuracy exceeding $99.79\%$. Even with precession and tidal effects, cAE maintains high precision across different $\Theta$ ranges. For efficiency, the cAE model also performs well. On GPU, it takes 0.1204 s to generate a single waveform, while IMRPhenomXP_NRTidalv2 on CPU takes 0.6663 s, giving a speed-up of about 5.54 times. As the batch size increases, this advantage becomes more clear. For batch size 1000, cAE–GPU takes 1.0210 s, while IMRPhenomXP_NRTidalv2 on CPU takes 10.1279 s, giving a speed-up of about 9.92 times. Residual-MLP–GPU shows similar results across all batch sizes. These results suggest that the cAE model maintains both high precision and relatively higher efficiency in waveform generation, especially when running on GPU.
Although this study has made some progress, several aspects remain worthy of further exploration. Future work can improve waveform reconstruction accuracy and generation efficiency by expanding the training dataset and incorporating more advanced model architectures. Generation speed is also a key factor, particularly critical for real-time gravitational wave applications. We plan to explore inference optimization strategies such as TensorRT [61, 62] and ONNX Runtime [63] to further reduce latency and enhance practical applicability. In addition, we will systematically evaluate the impact of latent variable dimensionality on waveform modeling and extend the parameter space to cover more complex physical systems, such as eccentric binaries and sources with stronger spin and tidal effects. On the architectural side, future work may explore diffusion models [64], Transformer variants [65], and variational autoencoders (VAEs, see FIG. 2(b)) to enhance generation performance and improve generalization.
@@ Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203004), the Fundamental Research Funds for the Central Universities Project (Grant No. 2024IAIS- ZD009), the National Natural Science Foundation of China (Grant No. 12575072, Grant No. 12347101), and the Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-MSX0103). The source code of this study is available at https://github.com/thisshy/BNS_Waveform_Generator_Based_on_CAE.
@@
Figure 13: Architectures of the residual-stacked Multilayer Perceptron (MLP) and Convolutional Neural Network (CNN) model.
Figure 14: Distribution of the $1\times 10^{6}$ training samples in the parameter space $\Theta$, including component masses, tidal deformabilities, and spin components. All parameters are sampled independently using uniform or isotropic priors.
Figure 15: Distribution of the $1\times 10^{5}$ testing samples. The parameter priors are identical to those used in training, ensuring a consistent and representative evaluation set.
Figure 16: Comparison of gravitational waveforms generated by the cAE (blue) and IMRPhenomXP_NRTidalv2 (yellow). The top subplot shows the full waveform over 022 s, with mismatch = $8.1\times 10^{-4}$ and 200 cycles. The middle subplot zooms into t=1.5998t=1.59981.99951.9995 s. The bottom subplot zooms into t=0t=00.59960.5996 s. The test waveform parameters are $m_{1}=2.84M_{\odot}$, $m_{2}=2.81M_{\odot}$, $\Lambda_{1}=188.48$, $\Lambda_{2}=233.67$, $\mathrm{spin}_{1x}=0.13$, $\mathrm{spin}_{1y}=-0.22$, $\mathrm{spin}_{1z}=-0.01$, $\mathrm{spin}_{2x}=-0.09$, $\mathrm{spin}_{2y}=-0.10$, $\mathrm{spin}_{2z}=-0.21$.
Figure 17: The top subplot shows the full waveform over 022 s, with mismatch = $1.2\times 10^{-3}$ and 250 cycles. The middle subplot zooms into t=1.5998t=1.59981.99951.9995 s. The bottom subplot zooms into t=0t=00.59960.5996 s. The test waveform parameters are $m_{1}=2.66M_{\odot}$, $m_{2}=1.81M_{\odot}$, $\Lambda_{1}=176.83$, $\Lambda_{2}=258.68$, $\mathrm{spin}_{1x}=0.19$, $\mathrm{spin}_{1y}=0.34$, $\mathrm{spin}_{1z}=0.08$, $\mathrm{spin}_{2x}=-0.02$, $\mathrm{spin}_{2y}=0.21$, $\mathrm{spin}_{2z}=-0.18$.
Figure 18: The top subplot shows the full waveform over 022 s, with mismatch = $3.4\times 10^{-3}$ and 300 cycles. The middle subplot zooms into t=1.5998t=1.59981.99951.9995 s. The bottom subplot zooms into t=0t=00.59960.5996 s. The test waveform parameters are $m_{1}=1.69M_{\odot}$, $m_{2}=1.66M_{\odot}$, $\Lambda_{1}=179.26$, $\Lambda_{2}=375.52$, $\mathrm{spin}_{1x}=-0.08$, $\mathrm{spin}_{1y}=0.20$, $\mathrm{spin}_{1z}=-0.36$, $\mathrm{spin}_{2x}=0.02$, $\mathrm{spin}_{2y}=0.36$, $\mathrm{spin}_{2z}=-0.04$.
Figure 19: The top subplot shows the full waveform over 022 s, with mismatch = $5.7\times 10^{-3}$ and 350 cycles. The middle subplot zooms into t=1.5998t=1.59981.99951.9995 s. The bottom subplot zooms into t=0t=00.59960.5996 s. The test waveform parameters are $m_{1}=1.31M_{\odot}$, $m_{2}=1.24M_{\odot}$, $\Lambda_{1}=135.05$, $\Lambda_{2}=417.58$, $\mathrm{spin}_{1x}=0.24$, $\mathrm{spin}_{1y}=0.03$, $\mathrm{spin}_{1z}=0.06$, $\mathrm{spin}_{2x}=-0.03$, $\mathrm{spin}_{2y}=-0.30$, $\mathrm{spin}_{2z}=0.32$.